rOA and rOB are the position vectors of the two particles Aand B at time t.If both particles start moving when t=0,determine whethert the particles collide and,if they do,give the value of t when this occurs and the position vector of the point of collision.If they do not ,find the time and their distance apart when they are closest together .

(a) rOA =3i-7j+(3i+2j)t, rOB=8i-6j+(2i+j)t.

(b) rOA=-5i+2j+9k+(5i-j+2k)t, rOB=i+2j+3k+(3i-j+4k)t.


1
Expert's answer
2021-05-07T09:31:31-0400

(a)

"\vec r_A=\vec r_B"

"3\vec i-7\vec j+(3\vec i+2\vec j)t=8\vec i-6\vec j+(2\vec i+\vec j)t"

"\begin{matrix}
 3+3t=8+2t \\
 -7+2t=-6+t \\

\end{matrix}"

"\begin{matrix}
 t=5 \\
 t=1 \\

\end{matrix}"

No solution. The particles will never collide.


"distance^2=d^2"

"=(8+2t-3-3t)^2+(-6+t+7-2t)^2"

"=(5-t)^2+(1-t)^2=2t^2-12t+26"

"t_{vertex}=\dfrac{12}{2(2)}=3"

"d_{min}=\sqrt{2(3)^2-12(3)+26}=2\sqrt{2}"

"\vec r_A(3)=12\vec i-\vec j"

"\vec r_B(3)=14\vec i-3\vec j"


(b)

"\vec r_A=\vec r_B"

"-5\vec i+2\vec j+9\vec k+(5\vec i-\vec j+2\vec k)t"

"=\vec i+2\vec j+3\vec k+(3\vec i-\vec j+4\vec k)t"

"\begin{matrix}
 -5+5t=1+3t \\
 2-t=2-t \\
9+2t=3+4t \\
\end{matrix}"

The solution "t=3"

The particles will collide at time "t=3" after the start, and


"\vec r_A=\vec r_B=10\vec i-\vec j+15\vec k"